OpenStudy (anonymous):

How many 2-digit numbers are there which are divisible by the sum of their digits ?

OpenStudy (anonymous):

Is it supposed to be an algorithm? Or your maths teacher is a kamikazi?

OpenStudy (anonymous):

no odd numbers have this property ( i guess!}

OpenStudy (anonymous):

I mean.. This is quite easy to do with a program but ffs trying it by hand its really hard

OpenStudy (anonymous):

Diogo - Its just the count of such two digit numbers :)

OpenStudy (anonymous):

i'i wrong - 63 is divisible by 9

OpenStudy (anonymous):

It's true for all products of 9.

OpenStudy (anonymous):

Except 99.

OpenStudy (anonymous):

right!

OpenStudy (anonymous):

yes indian, but first you need to find them all.. anyway, the sum of any number that is a multiple of 3 is divisible by 3. example: 15,21,24,27,30,33,36,39,42, etc

OpenStudy (anonymous):

99=3+(n-1)*3

OpenStudy (anonymous):

find n=??? that will your answer

OpenStudy (anonymous):

15 must be divisible by 1+5 = 6

OpenStudy (anonymous):

33

OpenStudy (anonymous):

ohhh. you're right mkuma36. I didnt understand the question at first. So 20 is divisible by 2, 30 by 3, 40 by 4, 50 by 5, 60 by 6, 70 by 7, 80 by 8, 90 by 9

OpenStudy (anonymous):

Exactly Diogo. But 33 is not the right answer I guess

OpenStudy (anonymous):

only 21 and 81 fit the bill for x1 up to 100

OpenStudy (anonymous):

yes Jimmyrep that is corect

OpenStudy (anonymous):

only 63 for x3 up to 100

OpenStudy (anonymous):

For x5, the only one that works is 50.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

this can be done by counting - not as tedious as i thought

OpenStudy (anonymous):

Even for the counting you can use math. Reduce the question into an eqn with two variables (the digits). It helps.

OpenStudy (anonymous):

33 also works for x3

OpenStudy (anonymous):

there are n x7 that works

OpenStudy (anonymous):

No dalvy.. 33 is not divisible by 6

OpenStudy (anonymous):

no x7

OpenStudy (anonymous):

I am silly.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

only 18 and 49 for x8

OpenStudy (anonymous):

lol 48!!!!!

OpenStudy (anonymous):

how it is 48??

OpenStudy (anonymous):

48 is divisible by 12.

OpenStudy (anonymous):

there are some weird results in number theory. for example the fraction of numbers between 1-100 which have two or more repetetive factors is 6 / pi^2! H ow the heck can pi be involved in something like that?

OpenStudy (anonymous):

Ok so I just did a visual basic program to solve this: here is the algorithm: Dim i, numero As Integer Dim j, a, k As Double For i = 10 To 99 j = i / 10 a = Math.Truncate(j) k = j - a k = k * 10 numero = k + a If i Mod numero = 0 Then MsgBox(i) End If Next ------------------------------------------- The values I found are: 10, 12, 18, 20, 21, 24, 27, 30, 36, 40, 42, 45, 48, 50, 54, 60, 63, 70, 72, 80, 81, 84, 90 Thats all. Good luck

OpenStudy (anonymous):

We know the divisibility rule of 3. It says that if the sum of digits are divisible by 3 or 9 then the number is divisible by 3 or 9. Can't be take the series as 3,6,9,12,15...

myininaya (myininaya):

i got 23 but it might wrong

myininaya (myininaya):

oh diogo got that many too i went a long process i wrote down every sum possible from 0 to 9 and check both digits and each sum to see if that number would be divisible by the sum and i ended up counting 23

OpenStudy (anonymous):

myininaya - you are correct and so is Diogo.. Good job mates. Diogo going out of the way to solve this problem... woohoo!!

myininaya (myininaya):

he actually has an algorithm up there

OpenStudy (anonymous):

Cool. Here is the short cut or so I think - (10a+b)/a+b=positive integer a+b+9a/a+b so we just hv 2 make 9a/a+b an positive integer value of A vary from 1 to 9 value of B vary from 0 to 9 so just put d values n u ll strt getting d reqd. numbers numbers are 10 12,21,24,42,48,84 20 multiples of 9=18,27,36,45,54,63,72,81 30 40 50 60 70 80 90

myininaya (myininaya):

thats cool

OpenStudy (anonymous):

myininaya, Diogo actually wrote a computer program to solve this.. Max respects!